rate of diffusion of co2 and so2 at same temp and pressure
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
As according to GRAHAM’s LAW; R1/R2 = \sqrt{M2/M1}for CO2; M1 = 44 So for SO2; M2 = 64 R1/R2 = \sqrt{64/44} = \sqrt{16/11}Hope it clears.........................................